Alycia opens her piggy bank to find only quarters and nickels inside. She counts the coins and finds that she has 57 coins totaling $8.65 How many quarters were in the piggy bank, and how many nickels were in there?
Found 2 solutions by ikleyn, Edwin McCravy:
Answer by ikleyn(52810)

.
Alycia opens her piggy bank to find only quarters and nickels inside. She counts the coins and finds that she has 57 coins totaling $8.65
How many quarters were in the piggy bank, and how many nickels were in there?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Let N = # of nickels.
The the number of quarters is (57-N).

The "value" equation is 

5N + 25*(57-N) = 865    cents.

5N + 1425 - 25N = 865,

-20N = 865 - 1425 = -560,

N = 560/20 = 28.

Answer. 28 nickels and 57 - 28 = 29 quarters.

There is entire bunch of lessons on coin problems
    - Coin problems
    - More Coin problems
    - Solving coin problems without using equations
    - Kevin and Randy Muise have a jar containing coins
    - Typical coin problems from the archive
    - More complicated coin problems
    - Solving coin problems mentally by grouping without using equations
    - Santa Claus helps solving coin problem
in this site.

Read them and become an expert in solution of coin problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Coin problems".


Answer by Edwin McCravy(20060)

First, without using algebra:

If all 57 coins had been nickels the total money would have 
been only 57×$0.05=$2.85.  But there was $8.65-$2.85=$5.80 
extra money that had to be accounted for by some of the coins 
being quarters which are worth 20 cents each more than a 
nickel. To find out how many quarters that had to be, we 
divide $5.80 by 20 cents or $0.20 and get 29 coins each of 
which were worth 20 cents more than a nickel, which were 
the quarters.  So there were 29 quarters and the other 
57-29=28 were nickels.

Second, by using algebra:

Coin equation:    q + n = 57
Money equation:  $0.25q + $0.05n = $8.65

Multiply the money equation by 100 and drop the $'s:

Money equation: 25q + 5n = 865

Divide money equation through by 5

Money equation: 5q + n = 173

Subtract the coin equation from the money equation:

                5q + n = 173
                 q + n =  57
                ------------
                4q     = 116
                     q =  29 quarters

                 q + n =  57
                29 + n =  57
                     n =  28 nickels.

Edwin